Given x1 = 3, x2 = 4, x3 = 8, x4 = -2, y1 = -6, y2 = 1, y3 = 5 and y4 = 0, find the value of the following:



1
Expert's answer
2021-05-24T18:56:12-0400

Given,

"x_1=3,x_2=4,x_3=8,x_4=-2\\y_1=-6,y_2=1,y_3=5,y_4=0"


So, "\sum_{i=1}^4 x_i y_i=\sum(x_1y_1+x_2y_2+x_3y_3+x_4y_4)"


"=(3)(-6)+(4)(1)+8(5)+(-2)(0)\\=-18+4+40\\=26"


Hence, "\sum_{i=1}^4 x_i y_i=26"